# Small-time asymptotics and the emergence of complex singularities for the KdV equation

### v1

@@ A.3 Matching into the inner region to determine $\Lambda$
$f_{0}\sim\sum_{n=0}^{\infty}\frac{a_{n}}{\xi^{3n+2}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,\quad a_{0}=A_{0}.$ (88)
@@
$\Lambda=\left(\frac{-2}{3^{3/2}}\right)^{3/2}\lim_{n\rightarrow\infty}\frac{(2/3)^{2n}a_{n}}{\Gamma(2n+{\textstyle\frac{3}{2}})}.$ (89)
It remains to consider the coefficients $a_{n}$ in order to take the limit in (89).
One way to determine the $a_{n}$ is to use the relationship between $f_{0}$ and $F$, given by (31), and then relying on (40) and (41), recalling that $\alpha$ and $A_{0}$ are related by (38). This approach gives
@@
where the $b_{n}$ are given by (41). Now, using the large $n$ behaviour of the $b_{n}$ in (42), we find
@@
Upon using the relationship between $\alpha$ and $A_{0}$ in (38), we find from (89) that $\Lambda$ is given by (17).
Another method is to substitute (88) into the governing equation (27) directly to give
@@
$a_{n}=(3n-1)(3n+1)a_{n-1}+\frac{2}{n}\sum_{m=0}^{n-1}(3m+2)a_{m}a_{n-m-1}.$ (90)
As a check, we have generated $a_{n}$ for large $n$ and then estimated the limit in (89) for various values of $A_{0}$ using repeated Shanks transformations. Combined with (89), this numerical exercise confirms (17).
A.4 The special case $A_{0}=-N(N+1)$
For an initial condition with double poles whose local behaviour is given by (3), our leading-order prediction (18) with (17) suggests that the amplitude of the dispersive waves vanishes for triangular values $A_{0}=-N(N+1)$. However, this absence of dispersive waves for certain values of $A_{0}$ is known to hold only for the exceptional $\mathrm{sech}^{2}$-type initial conditions (4). Otherwise, when $A_{0}=-N(N+1)$ we need to revisit our analysis and consider higher-order contributions, the details of which we summarise here.
To proceed, we will need some results from appendix B below. First, we will need to consider the first two terms in (21), which are characterised by $A_{0}$ and the residue $A_{1}$. We shall also need the first two terms in the expansion for $f(\xi)$, given by (97), where now the correction term $f_{1}$ satisfies the ode (98). Using the far-field behaviour (100)–(101), we can integrate the equation for $f_{1}$ to give the second-order linear equation (120). For our purposes, we shall write out the full series for $f_{1}$ in (101) as
$f_{1}\sim A_{1}\sum_{n=0}^{\infty}\frac{c_{n}}{\xi^{3n+1}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (91)
@@ A.4 The special case $A_{0}=-N(N+1)$
A key point is that, in general, both of the series in (100) and (101) are divergent; however, for the special case $A_{0}=-N(N+1)$ the series (100) converges (for sufficiently large $|\xi|$) to the appropriate rational solution described in subsection 3.4 (the first four of which are listed in (52)–(55)). Therefore, when $A_{0}=-N(N+1)$, the crucial divergence that is driving the exponentially-small terms (that are turned on across Stokes lines) comes from $f_{1}$ and not $f_{0}$. Thus, our inner expansion for $A_{0}=-N(N+1)$ can be interpreted as being
$u(x,t)\sim\frac{1}{(3t)^{2/3}}\,\left[\mbox{rational soln}\,f_{0}(\xi)\right]+\frac{1}{(3t)^{1/3}}f_{1}(\xi)\quad\mbox{as}\quad t\rightarrow 0^{+},$ (92)
where $f_{1}$ has the expansion (91).
Returning to our outer expansion, we found that our late-order terms are of the form (15). To be consistent with (92), we must now choose $\gamma=1/2$ (and not $\gamma=3/2$), in which case the late-order terms behave as
@@
as $n\rightarrow\infty$ (and not (87)). Matching with the inner region (91)–(92), we find
$\Lambda=\left(\frac{-2}{3^{3/2}}\right)^{1/2}A_{1}\lim_{n\rightarrow\infty}\frac{(2/3)^{2n}c_{n}}{\Gamma(2n+{\textstyle\frac{1}{2}})}.$ (93)
@@
We can obtain a recurrence relation for $c_{n}$ by substituting (91) into the governing ode for $f_{1}$, namely (120) (which we formulate below in appendix B), to give
@@
Thus, for $A_{0}=-2$ we have from (93) that
@@
In summary, for the special cases $A_{0}=-N(N+1)$, where $N\in\mathbb{N}$, instead of the exponentially-small contribution (16) being switched on across a Stokes line $\mathrm{arg}(x-x_{0})=-2\pi/3$, we have
@@
Together with the analogous contribution from the lower-half plane, the result is that, on the real line, instead of (18), we have
@@
as $x\rightarrow$-$\infty$, $t\rightarrow 0^{+}$. Compared with the generic case (5), the dispersive waves are algebraically smaller by a scaling of $(-x/t)^{1/2}$ due to the term $(-1/xt)^{1/4}$ compared to $(-x/t^{3})^{1/4}$.
@@ Appendix B Higher-order terms in matched asymptotic expansions and special cases
In this appendix, we consider higher-order terms in our expansions in order to demonstrate how further terms in (29) could be generated. Further, we use these results to clarify a couple of special cases, including $A_{1}=0$ and $A_{0}=-2$.
@@ B.1 Outer and inner expansions with more terms
Here we shall consider more terms in our expansions, starting with our initial condition and then our inner problem. As such, we shall revisit the outer expansion, but this time including more details. To progress, we extend the expansion for our class of initial conditions (3) to be (21). The first three terms of the expansion (10) are
$u\sim u_{0}(x)+tu_{1}(x)+t^{2}u_{2}(x)\quad\mbox{as}\quad t\rightarrow 0^{+},$ (94)
@@
$u_{1}\sim\frac{12A_{0}(A_{0}+2)}{(x-x_{0})^{5}}+\frac{6A_{1}(3A_{0}+1)}{(x-x_{0})^{4}}+\frac{6(A_{1}^{2}+2A_{0}A_{2})}{(x-x_{0})^{3}}\quad\mbox{as}\quad x\rightarrow x_{0},$ (95)
$u_{2}\sim\frac{252A_{0}(A_{0}+2)(A_{0}+5)}{(x-x_{0})^{8}}+\frac{180A_{1}(3A_{0}^{2}+9A_{0}+2)}{(x-x_{0})^{7}}\quad\mbox{as}\quad x\rightarrow x_{0}.$ (96)
We now combine (94) and (96) to give
@@
Therefore, for the inner problem (25) with pde (26), we can write
$f\sim f_{0}(\xi)+(3t)^{1/3}f_{1}(\xi)+(3t)^{2/3}f_{2}(\xi)\quad\mbox{as}\quad t\rightarrow 0^{+},$ (97)
so that $f_{0}$ satisfies (27), while $f_{1}$ and $f_{2}$ satisfy
$f_{1}-\xi f_{1}^{\prime}+6(f_{0}f_{1}^{\prime}+f_{0}^{\prime}f_{1})+f_{1}^{\prime\prime\prime}=0,$ (98)
$\xi f_{2}^{\prime}+6(f_{0}f_{2}^{\prime}+f_{1}f_{1}^{\prime}+f_{0}^{\prime}f_{2})+f_{2}^{\prime\prime\prime}=0,$ (99)
@@
$f_{0}$ $\sim\frac{A_{0}}{\xi^{2}}+\frac{4A_{0}(A_{0}+2)}{\xi^{5}}+\frac{28A_{0}(A_{0}+2)(A_{0}+5)}{\xi^{8}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (100)
$f_{1}$ $\sim\frac{A_{1}}{\xi}+\frac{2A_{1}(3A_{0}+1)}{\xi^{4}}+\frac{20A_{1}(3A_{0}^{2}+9A_{0}+2)}{\xi^{7}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (101)
$f_{2}$ $\sim A_{2}+\frac{2(A_{1}^{2}+2A_{0}A_{2})}{\xi^{3}}+\frac{40A_{0}A_{2}(A_{0}+2)+10A_{1}^{2}(4A_{0}+3)}{\xi^{6}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (102)
together with other conditions suppressing exponentially growing correction terms, which we will discuss below where appropriate. (Note that the terms in (100)–(102) could also be determined directly from (27), (98) or (99); for example, (100) contains the first few terms of (88) together with the recurrence relation (90).)
In general, the solutions for $f_{0}$ will be related to our decreasing tritronquée solutions $F$ of PII via (31). Thus, $f_{0}$ will generically have an infinite number of double poles at points that we label $\xi=\xi_{0}$. As we have already noted in (29), this means that for each $\xi=\xi_{0}$, there is a corresponding singularity of our KdV solution that behaves like $s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}$ as $t\rightarrow 0^{+}$.
By considering (27), we find $f_{0}$ has the local behaviour about each of its singularities of the form
$f_{0}\sim\frac{-2}{(\xi-\xi_{0})^{2}}+{\textstyle\frac{1}{6}}\xi_{0}+\mathcal{O}\left((\xi-\xi_{0})^{2}\right)\quad\mbox{as}\quad\xi\rightarrow\xi_{0}.$ (103)
Turning to the problem for $f_{1}$, we can integrate (98) together with (101) to give
$f_{1}^{\prime\prime}+(6f_{0}-\xi)f_{1}+A_{1}=0.$ (104)
As $f_{1}$ satisfies a linear equation, it will also be singular at each of the points $\xi=\xi_{0}$. Regardless of the precise form of the solution for $f_{1}$, the governing equation (120) can be used to show $f_{1}$ will behave as
$f_{1}\sim\frac{\mu_{1}}{(\xi-\xi_{0})^{3}}-{\textstyle\frac{1}{12}}\mu_{1}+\mathcal{O}(\xi-\xi_{0})\quad\mbox{as}\quad\xi\rightarrow\xi_{0},$ (105)
where $\mu_{1}$ is some constant that depends on $A_{1}$ (to compute $\mu_{1}$, the full solution $f_{1}$ would need to be calculated). Finally, directly from (99) we find
$f_{2}\sim-\frac{3\mu_{1}^{2}}{8(\xi-\xi_{0})^{4}}+\frac{\mu_{2}}{(\xi-\xi_{0})^{3}}\quad\mbox{as}\quad\xi\rightarrow\xi_{0},$ (106)
@@ B.2 Higher-order terms for singularity location $x=s(t)$
In order to derive further terms in our expansion for $x=s(t)$, the location of each singularity of our KdV solution, we first note that including the next nonzero term in the expansion (2) gives
$u\sim-\frac{2}{(x-s(t))^{2}}+\frac{\dot{s}}{6}\quad\mbox{as}\quad x\rightarrow s(t).$ (107)
@@
$s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}+(3t)^{2/3}\xi_{1}+(3t)\xi_{2}\quad\mbox{as}\quad t\rightarrow 0^{+},$ (108)
@@
$\left.+(3t)^{2/3}\left(-\frac{6\xi_{1}^{2}}{(\xi-\xi_{0})^{4}}-\frac{4\xi_{2}}{(\xi-\xi_{0})^{3}}+\ldots\right)+\ldots\right]\quad\mbox{as}\quad x\rightarrow x_{0},\quad t\rightarrow 0^{+},$ (109)
where we have been careful not to include terms in (109) that are neglected in (107).
Now, given the expansion (97), we can infer from (109) that
$f_{0}\sim$ $\frac{-2}{(\xi-\xi_{0})^{2}}+\frac{1}{6}\xi_{0}+\ldots,$ (110)
$f_{1}\sim$ $-\frac{4\xi_{1}}{(\xi-\xi_{0})^{3}}+\frac{1}{3}\xi_{1}+\ldots,$ (111)
$f_{2}\sim$ $-\frac{6\xi_{1}^{2}}{(\xi-\xi_{0})^{4}}-\frac{4\xi_{2}}{(\xi-\xi_{0})^{3}}+\ldots\quad\mbox{as}\quad x\rightarrow x_{0}.$ (112)
Thus, by comparing with (105)–(106), we find
$\xi_{1}=-{\textstyle\frac{1}{4}}\mu_{1},\quad\xi_{2}=-{\textstyle\frac{1}{4}}\mu_{2}.$ (113)
@@
$s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}-{\textstyle\frac{1}{4}}(3t)^{2/3}\mu_{1}-{\textstyle\frac{1}{4}}(3t)\mu_{2}\quad\mbox{as}\quad t\rightarrow 0^{+},$ (114)

### v2

@@ A.3 Matching in inner region to determine $\Lambda$
$f_{0}\sim\sum_{n=0}^{\infty}\frac{a_{n}}{\xi^{3n+2}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,\quad a_{0}=A_{0}.$ (A.5)
@@
$\Lambda=\left(\frac{-2}{3^{3/2}}\right)^{3/2}\lim_{n\rightarrow\infty}\frac{(2/3)^{2n}a_{n}}{\Gamma(2n+{\textstyle\frac{3}{2}})}.$ (A.6)
It remains to consider the coefficients $a_{n}$ in order to take the limit in (A.6).
One way to determine the $a_{n}$ is to use the relationship between $f_{0}$ and $F$, given by (3.9), and then relying on (3.20) and (3.21), recalling that $\alpha$ and $A_{0}$ are related by (3.16). This approach gives
@@
where the $b_{n}$ are given by (3.21). Now, using the large $n$ behaviour of the $b_{n}$ in (3.22), we find
@@
Upon using the relationship between $\alpha$ and $A_{0}$ in (3.16), we find from (A.6) that $\Lambda$ is given by (2.11).
Another method is to substitute (A.5) into the governing equation (3.5) directly to give
@@
$a_{n}=(3n-1)(3n+1)a_{n-1}+\frac{2}{n}\sum_{m=0}^{n-1}(3m+2)a_{m}a_{n-m-1}.$ (A.7)
As a check, we have generated $a_{n}$ for large $n$ and then estimated the limit in (A.6) for various values of $A_{0}$ using repeated Shanks transformations. Combined with (A.6), this numerical exercise confirms (2.11).
A.4 Special case $A_{0}=-N(N+1)$
For an initial condition with double poles whose local behaviour is given by (1.3), our leading-order prediction (2.12) with (2.11) suggests that the amplitude of the dispersive waves vanishes for triangular values $A_{0}=-N(N+1)$. However, this absence of dispersive waves for certain values of $A_{0}$ is known to hold only for the exceptional $\mathrm{sech}^{2}$-type initial conditions (1.4). Otherwise, when $A_{0}=-N(N+1)$ we need to revisit our analysis and consider higher-order contributions, the details of which we summarise here.
To proceed, we will need some results from appendix B.1 below. First, we will need to consider the first two terms in (2.15), which are characterised by $A_{0}$ and the residue $A_{1}$. We shall also need the first two terms in the expansion for $f(\xi)$, given by (B.4), where now the correction term $f_{1}$ satisfies the ode (B.5). Using the far-field behaviour (B.7)–(B.8), we can integrate the equation for $f_{1}$ to give the second-order linear equation (B.27). For our purposes, we shall write out the full series for $f_{1}$ in (B.8) as
$f_{1}\sim A_{1}\sum_{n=0}^{\infty}\frac{c_{n}}{\xi^{3n+1}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (A.8)
@@ A.4 Special case $A_{0}=-N(N+1)$
A key point is that, in general, both of the series in (B.7) and (B.8) are divergent; however, for the special case $A_{0}=-N(N+1)$ the series (B.7) converges (for sufficiently large $|\xi|$) to the appropriate rational solution described in subsection 3.5 (the first four of which are listed in (3.33)–(3.36)). Therefore, when $A_{0}=-N(N+1)$, the crucial divergence that is driving the exponentially small terms (that are turned on across Stokes lines) comes from $f_{1}$ and not $f_{0}$. Thus, our inner expansion for $A_{0}=-N(N+1)$ can be interpreted as being
$u(x,t)\sim\frac{1}{(3t)^{2/3}}\,\left[\mbox{rational soln}\,f_{0}(\xi)\right]+\frac{1}{(3t)^{1/3}}f_{1}(\xi)\quad\mbox{as}\quad t\rightarrow 0^{+},$ (A.9)
where $f_{1}$ has the expansion (A.8).
Returning to our outer expansion, we found that our late-order terms are of the form (2.9). To be consistent with (A.9), we must now choose $\gamma=1/2$ (and not $\gamma=3/2$), in which case the late-order terms behave as
@@
as $n\rightarrow\infty$ (and not (A.4)). Matching with the inner region (A.8)–(A.9), we find
$\Lambda=\left(\frac{-2}{3^{3/2}}\right)^{1/2}A_{1}\lim_{n\rightarrow\infty}\frac{(2/3)^{2n}c_{n}}{\Gamma(2n+{\textstyle\frac{1}{2}})}.$ (A.10)
@@
We can obtain a recurrence relation for $c_{n}$ by substituting (A.8) into (B.27)
@@
Thus, for $A_{0}=-2$ we have from (A.10) that
@@
In summary, for the special cases $A_{0}=-N(N+1)$, where $N\in\mathbb{N}$, instead of the exponentially small contribution (2.10) being switched on across a Stokes line $\mathrm{arg}(x-x_{0})=-2\pi/3$, we have
@@
Together with the analogous contribution from the lower-half plane, the result is that, on the real line, instead of (2.12), we have
@@
as $x\rightarrow$-$\infty$, $t\rightarrow 0^{+}$. Compared with the generic case (1.5), the dispersive waves are algebraically smaller by a scaling of $(-x/t)^{1/2}$ due to the term $(-1/xt)^{1/4}$ compared to $(-x/t^{3})^{1/4}$.
@@ Appendix B Higher-order terms in matched asymptotic expansions and special cases
In this appendix, we consider higher-order terms in our expansions in order to demonstrate how further terms in (3.7) could be generated. Further, we use these results to clarify a couple of special cases, including $A_{1}=0$ and $A_{0}=-2$.
@@ B.1 Outer and inner expansions with more terms
Here we shall consider more terms in our expansions, starting with our initial condition and then our inner problem. As such, we shall revisit the outer expansion, but this time including more details. To progress, we extend the expansion for our class of initial conditions (1.3) to be (2.15). The first three terms of the expansion (2.4) are
$u\sim u_{0}(x)+tu_{1}(x)+t^{2}u_{2}(x)\quad\mbox{as}\quad t\rightarrow 0^{+},$ (B.1)
@@
$u_{1}\sim\frac{12A_{0}(A_{0}+2)}{(x-x_{0})^{5}}+\frac{6A_{1}(3A_{0}+1)}{(x-x_{0})^{4}}+\frac{6(A_{1}^{2}+2A_{0}A_{2})}{(x-x_{0})^{3}}\quad\mbox{as}\quad x\rightarrow x_{0},$ (B.2)
$u_{2}\sim\frac{252A_{0}(A_{0}+2)(A_{0}+5)}{(x-x_{0})^{8}}+\frac{180A_{1}(3A_{0}^{2}+9A_{0}+2)}{(x-x_{0})^{7}}\quad\mbox{as}\quad x\rightarrow x_{0}.$ (B.3)
We now combine (B.1) and (B.3) to give
@@
Therefore, for the inner problem (3.3) with pde (3.4), we can write
$f\sim f_{0}(\xi)+(3t)^{1/3}f_{1}(\xi)+(3t)^{2/3}f_{2}(\xi)\quad\mbox{as}\quad t\rightarrow 0^{+},$ (B.4)
so that $f_{0}$ satisfies (3.5), while $f_{1}$ and $f_{2}$ satisfy
$f_{1}-\xi\frac{\mathrm{d}f_{1}}{\mathrm{d}\xi}+6\left(f_{0}\frac{\mathrm{d}f_{1}}{\mathrm{d}\xi}+\frac{\mathrm{d}f_{0}}{\mathrm{d}\xi}f_{1}\right)+\frac{\mathrm{d}^{3}f_{1}}{\mathrm{d}\xi^{3}}=0,$ (B.5)
$\xi\frac{\mathrm{d}f_{2}}{\mathrm{d}\xi}+6\left(f_{0}\frac{\mathrm{d}f_{2}}{\mathrm{d}\xi}+f_{1}\frac{\mathrm{d}f_{1}}{\mathrm{d}\xi}+\frac{\mathrm{d}f_{0}}{\mathrm{d}\xi}f_{2}\right)+\frac{\mathrm{d}^{3}f_{2}}{\mathrm{d}\xi^{3}}=0,$ (B.6)
@@
$f_{0}$ $\sim\frac{A_{0}}{\xi^{2}}+\frac{4A_{0}(A_{0}+2)}{\xi^{5}}+\frac{28A_{0}(A_{0}+2)(A_{0}+5)}{\xi^{8}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (B.7)
$f_{1}$ $\sim\frac{A_{1}}{\xi}+\frac{2A_{1}(3A_{0}+1)}{\xi^{4}}+\frac{20A_{1}(3A_{0}^{2}+9A_{0}+2)}{\xi^{7}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (B.8)
$f_{2}$ $\sim A_{2}+\frac{2(A_{1}^{2}+2A_{0}A_{2})}{\xi^{3}}+\frac{40A_{0}A_{2}(A_{0}+2)+10A_{1}^{2}(4A_{0}+3)}{\xi^{6}}\quad\mbox{as}\quad\xi\rightarrow-\mathrm{i}\infty,$ (B.9)
together with other conditions suppressing exponentially growing correction terms, which we will discuss below where appropriate. (Note that the terms in (B.7)–(B.9) could also be determined directly from (3.5), (B.5) or (B.6); for example, (B.7) contains the first few terms of (A.5) together with the recurrence relation (A.7).)
In general, the solutions for $f_{0}$ will be related to our decreasing tritronquée solutions $F$ of PII via (3.9). Thus, $f_{0}$ will generically have an infinite number of double poles at points that we label $\xi=\xi_{0}$. As we have already noted in (3.7), this means that for each $\xi=\xi_{0}$, there is a corresponding singularity of our KdV solution that behaves like $s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}$ as $t\rightarrow 0^{+}$.
By considering (3.5), we find $f_{0}$ has the local behaviour about each of its singularities of the form
$f_{0}\sim\frac{-2}{(\xi-\xi_{0})^{2}}+{\textstyle\frac{1}{6}}\xi_{0}+\mathcal{O}\left((\xi-\xi_{0})^{2}\right)\quad\mbox{as}\quad\xi\rightarrow\xi_{0}.$ (B.10)
Turning to the problem for $f_{1}$, we can integrate (B.5) together with (B.8) to give
$\frac{\mathrm{d}^{2}f_{1}}{\mathrm{d}\xi^{2}}+(6f_{0}-\xi)f_{1}+A_{1}=0.$ (B.11)
As $f_{1}$ satisfies a linear equation, it will also be singular at each of the points $\xi=\xi_{0}$. Regardless of the precise form of the solution for $f_{1}$, the governing equation (B.27) can be used to show $f_{1}$ will behave as
$f_{1}\sim\frac{\mu_{1}}{(\xi-\xi_{0})^{3}}-{\textstyle\frac{1}{12}}\mu_{1}+\mathcal{O}(\xi-\xi_{0})\quad\mbox{as}\quad\xi\rightarrow\xi_{0},$ (B.12)
where $\mu_{1}$ is some constant that depends on $A_{1}$ (to compute $\mu_{1}$, the full solution $f_{1}$ would need to be calculated). Finally, directly from (B.6) we find
$f_{2}\sim-\frac{3\mu_{1}^{2}}{8(\xi-\xi_{0})^{4}}+\frac{\mu_{2}}{(\xi-\xi_{0})^{3}}\quad\mbox{as}\quad\xi\rightarrow\xi_{0},$ (B.13)
@@ B.2 Higher-order terms for singularity location $x=s(t)$
In order to derive further terms in our expansion for $x=s(t)$, the location of each singularity of our KdV solution, we first note that including the next nonzero term in the expansion (1.2) gives
$u\sim-\frac{2}{(x-s(t))^{2}}+\frac{\dot{s}}{6}\quad\mbox{as}\quad x\rightarrow s(t).$ (B.14)
@@
$s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}+(3t)^{2/3}\xi_{1}+(3t)\xi_{2}\quad\mbox{as}\quad t\rightarrow 0^{+},$ (B.15)
@@
$\left.+(3t)^{2/3}\left(-\frac{6\xi_{1}^{2}}{(\xi-\xi_{0})^{4}}-\frac{4\xi_{2}}{(\xi-\xi_{0})^{3}}+\ldots\right)+\ldots\right]\quad\mbox{as}\quad x\rightarrow x_{0},\quad t\rightarrow 0^{+},$ (B.16)
where we have been careful not to include terms in (B.16) that are neglected in (B.14).
Now, given the expansion (B.4), we can infer from (B.16) that
$f_{0}\sim$ $\frac{-2}{(\xi-\xi_{0})^{2}}+\frac{1}{6}\xi_{0}+\ldots,$ (B.17)
$f_{1}\sim$ $-\frac{4\xi_{1}}{(\xi-\xi_{0})^{3}}+\frac{1}{3}\xi_{1}+\ldots,$ (B.18)
$f_{2}\sim$ $-\frac{6\xi_{1}^{2}}{(\xi-\xi_{0})^{4}}-\frac{4\xi_{2}}{(\xi-\xi_{0})^{3}}+\ldots\quad\mbox{as}\quad x\rightarrow x_{0}.$ (B.19)
Thus, by comparing with (B.12)–(B.13), we find
$\xi_{1}=-{\textstyle\frac{1}{4}}\mu_{1},\quad\xi_{2}=-{\textstyle\frac{1}{4}}\mu_{2}.$ (B.20)
@@
$s(t)\sim x_{0}+(3t)^{1/3}\xi_{0}-{\textstyle\frac{1}{4}}(3t)^{2/3}\mu_{1}-{\textstyle\frac{1}{4}}(3t)\mu_{2}\quad\mbox{as}\quad t\rightarrow 0^{+},$ (B.21)